OpenStudy (anonymous):

What is a reasonable estimate for the area of a circle with radius 4 cm? 12 cm2 16 cm2 24 cm2 48 cm2

OpenStudy (nurali):

r = 4cm Area = πr^2 = 22/7 * 4^2 cm^2 = 352/7 cm^2 = 50.29 cm^2 Therefore, Answer is D. A =48 cm^2

OpenStudy (anonymous):

\[A=\pi r ^{2}=\pi4^{2}=16\pi \] substitute the value of pi and get the solution.

OpenStudy (anonymous):

ok thanks

OpenStudy (nurali):

Anytime.